### v1

# RouterArena: An Open Platform for Comprehensive Comparison of LLM Routers

Abstract
Today’s LLM ecosystem comprises a wide spectrum of models that differ in size, capability, and cost. No single model is optimal for all scenarios; hence, LLM routers have become essential for selecting the most appropriate model under varying circumstances. However, the rapid emergence of various routers makes choosing the right one increasingly challenging. To address this problem, we need comprehensive router comparison and a standardized leaderboard, similar to those available for models. In this work, we introduce RouterArena, the first open platform enabling comprehensive comparison of LLM routers. RouterArena has (1) a principally constructed dataset with broad knowledge domain coverage, (2) distinguishable difficulty levels for each domain, (3) an extensive list of evaluation metrics, and (4) an automated framework for leaderboard updates. Leveraging our framework, we have produced the initial leaderboard with detailed metrics comparison as shown in Figure 1. We will soon open our platform to the public.
Figure 1: A quick view of RouterArena leaderboard and performance-cost trade-off.
1 Introduction
Large Language Models (LLMs) are rapidly diversifying, offering an ever-wider spectrum of capabilities and inference costs. This diversity increasingly challenges the prevailing LLM usage pattern in which users manually choose models for their queries. The difficulty stems from the fact that no single model is universally optimal: powerful models excel at complex tasks but are costly, while smaller models are more efficient yet may struggle on difficult queries. As a result, LLM routers that automatically select models based on input queries are increasingly recognized as a core system primitive in practical deployments.
Given its importance and promise, many LLM routers have recently emerged in both industry and academia (Figure 2). A notable example is GPT-5 [27], which incorporates routing as a key feature by directing user queries to the most suitable model within the OpenAI family. As routers proliferate, the challenge shifts from selecting the right model to selecting the right router. Unfortunately, router evaluation has not kept pace: there is currently no open evaluation platform, akin to LMArena [5], that systematically compares open-source routers [15, 42] and commercial routing services [25, 24] under a unified protocol.
It is urgent to fill this gap by building a Router Arena that can comprehensively evaluate and rank routers, enabling users to understand the status quo and make informed choices. However, unlike model arenas, designing a router arena is considerably more challenging due to the requirements from three key aspects. (1) Dataset. To evaluate whether a router can recognize problem domains and dispatch queries to appropriate models at minimal cost, the arena dataset must cover a broad range of domains and subjects, as well as varying difficulty levels. (2) Metrics. Router performance is inherently multi-dimensional, and so should be the arena ranking. While accuracy and cost are the primary metrics, it is also important to capture other dimensions such as routing optimality and robustness. (3) Framework. To enable live leaderboard updates, the arena must have a user-friendly framework that can automatically evaluate new open-source and commercial routers. Although existing studies explore some of these directions, as summarized in Table 1 and discussed in Section 2, they fail to address each challenge in a comprehensive way.
In this work, we present RouterArena, the first open platform for comprehensive evaluation and comparison of LLM routers. It addresses the above key challenges with the following designs:
  • A Principled Diverse Dataset. To ensure broad coverage, we construct the dataset using the Dewey Decimal Classification system adopted in libraries, covering all domains except religion. For each subject, we apply Bloom’s taxonomy to design queries at three difficulty levels, producing a diverse dataset of $\sim$8,000 queries spanning 9 domains and 44 categories for router evaluation.
  • Extensive Metrics for Arena Ranking. We construct router leaderboards by considering an extensive list of deployment-relevant metrics including query-answer accuracy, query-answer cost, routing optimality (cheapest correct selection), robustness to query perturbations (consistency), and router overhead (latency). This enables router comparison from multiple perspectives.
  • An Automated Framework for Leaderboard Updates. We design a framework that automatically evaluates new routers, collects metrics, and updates the leaderboard. The framework supports both open-source and commercial routers, and employs prefix caching to improve efficiency.
Figure 1 provides a quick view of our accuracy–cost leaderboard along with other details. We have found that although GPT-5 achieves higher accuracy, its cost is significantly higher than that of other routers due to its model pool being restricted to the OpenAI family. Consequently, it does not rank as the best router on our accuracy–cost leaderboard.
Our vision is for RouterArena to serve as an open community venue for evaluating routers as the ecosystem evolves, providing a standardized basis for fair comparison and progress tracking. By lowering the barrier to evaluation and enabling transparent, reproducible results, RouterArena will help researchers and practitioners design, improve, and adopt better routers.
2 Motivation
The Rapid Emergence of LLM Routers.
As shown in Figure 2, the landscape of LLM routers is rapidly expanding, evolving from academic exploration to commercial deployment. From a few scattered academia routers [39, 2, 13] in mid-2023, the number of publications expanded to more than a dozen by 2024 [22, 43, 3, 41]. By 2025, not only did academia routers continue to grow [36, 16, 9, 40], but commercial products also emerged [25, 24], most notably GPT-5 [27] with a built-in router.
New Problem: How to Choose the Right Router?
Figure 2: Timeline of example router-related works and products.
Table 1: Comparison of existing works [15, 17, 11, 42] and RouterArena. RouterArena enables comprehensive router comparison with extensive query categories, difficulty levels, evaluation metrics, and router inclusion.
Benchmark Query categories Difficulty levels Evaluation Metrics Commercial Routers Router Ranking
RouterBench 24 Categories ✗ No analysis ✗ Deferral curve only
RouterEval 27 Categories ✗ No analysis ✗ Accuracy metric only
FusionBench 26 Categories LLM-judge analysis ✗ Deferral curve only
EmbedLLM 26 Categories ✗ No analysis ✗ Accuracy metric only
RouterArena
44 Categories
based on DDC
✓ 5 Bloom Level
Classification
✓ 5 Evaluation
perspectives
✓ 3 Included
✓ Multi-metric
leaderboard
Answering this question requires comprehensive router comparisons to understand the current landscape. Such comprehensiveness entails dataset categories, difficulty levels, evaluation metrics, and router inclusion. However, our review of existing work reveals that no such comprehensive comparisons are available today. As shown in Table 1, the existing work falls short in the following aspects.
  • Narrow Query Category Coverage. They lack full coverage of query categories, making them impossible to evaluate router performance on queries from excluded categories.
  • Indistinguishable Difficulty Levels. They do not differentiate queries by difficulty, limiting their ability to test accuracy–cost tradeoffs.
  • Incomplete Evaluation Metrics. They only consider a subset of relevant metrics, overlooking important dimensions such as optimality, robustness, and latency.
  • No Support of Commercial Routers. Current frameworks evaluate only open-source routers and do not extend to closed-source or commercial routers.
  • No Router Leaderboard. There is no leaderboard that allows people to compare all routers under a unified evaluation protocol.
This Work: RouterArena.
This gap motivates us to design RouterArena, an open platform for comprehensive router comparisons. In the remainder of this paper, we first introduce the key components of RouterArena: principled dataset construction, comprehensive metric formulation, and an automated evaluation framework with live leaderboard. We then present our evaluation results and discuss the key findings.
3 RouterArena Evaluation Dataset
To enable meaningful and unbiased router comparisons, a high-quality evaluation dataset is essential. In this work, we construct such a dataset by adhering to two guiding principles for data collection.
Principle 1: DDC-Inspired Diverse Domain Coverage.
To evaluate a router’s ability to recognize problem domains and route queries to the appropriate specialist models, the dataset must provide broad domain coverage. To achieve this, we draw inspiration from the Dewey Decimal Classification (DDC) system [8], a book classification framework widely used in libraries. The DDC is renowned for its comprehensive and logical structure, providing a proven methodology for organizing the entire world of knowledge into distinct, hierarchical categories [30].
Principle 2: Bloom-Based Distinguishable Difficulty Levels.
To evaluate whether a router can determine query difficulty and make accuracy-cost tradeoffs—choosing between powerful but expensive models and weaker but cheaper ones—the dataset must include clearly distinguishable difficulty levels. To structure these difficulty levels, we adopt Bloom’s taxonomy [1], a widely used framework in quantifying question complexity [34, 14, 28]. It works by classifying cognitive tasks into six ascending categories: remembering, understanding, applying, analyzing, evaluating, and creating. In this work, we further group these six levels into easy-medium-difficult three levels.
Dataset Construction Process.
Following these two principles, we curate our evaluation dataset as follows. To ensure coverage across all DDC categories (excluding religion), we first collect all the queries from two existing LLM benchmark datasets and then supplement underrepresented categories with data from 21 open-source, domain-specific datasets. To determine the difficulty level of each query based on Bloom’s taxonomy, we employ an LLM-as-Judge approach with DeepSeek-V3.1 [7] (prompt specified in Appendix D), enabling automatic difficulty annotation. We exclude the create-type questions because they are open-ended and cannot be reliably evaluated. We then categorize remembering and understanding questions as easy, applying questions as medium, and analyzing and evaluating questions as difficult.
Next, to fairly distribute questions across categories and difficulty levels, we propose a recursive deficit redistribution algorithm. We begin by setting the ratio of science to humanities at 2:1. Within each top-level category, if a sub-category falls short of its proportional quota, the resulting surplus is recursively and uniformly redistributed to those sub-categories that exceed their initial allocation. We apply the same procedure within each sub-category to allocate data across different difficulty levels, ensuring balanced coverage throughout the dataset.
The above steps yield approximately 62,000 queries in total. However, this raw dataset contains many highly similar or even duplicate questions inherited from the various sampled sources. Such redundancy does not benefit router evaluation and may even introduce noise into the results. To address this, we perform cosine-similarity–based de-duplication using sentence-transformers/all-MiniLM-L6-v2. By strictly following the allocation strategy and selecting the least similar samples, we ensure that the resulting RouterArena dataset maintains broad coverage with minimal redundancy.
The Resulting Dataset.
Our final evaluation dataset consists of 8,400 queries sampled from 23 source datasets. It spans nine top-level domains and 44 categories, with each category containing queries across three difficulty levels. Figure 3 illustrates the detailed composition of the dataset. Note that the distribution of difficulty levels is skewed, but it reflects real-world query patterns—easy and medium questions occurring more frequently than hard ones. We include more details about the dataset in Appendix B.1, including dataset schema and concrete query examples.
Figure 3: Dataset composition. For ease of demonstration, we merged some categories.
4 RouterArena Evaluation Metrics
RouterArena supports comprehensive router evaluation along five dimensions.
(1) Query-answer Accuracy.
This metric captures a router’s ability to direct queries to the appropriate models such that they are correctly answered. We calculate accuracy as the average correctness across all our dataset queries.
(2) Query-answer Cost.
This measures the cost incurred by a router’s routing decisions. To address important factors such as the variable cost introduced by input length and generation length (e.g., in chain-of-thought reasoning) as well as the distinct computational characteristics of Mixture-of-Experts (MoE) models, we use the actual inference cost measured by:
Where $c$ is the cost per token and $N$ is the number of tokens. We obtain the cost $c$ for the specific models a router chooses using the official API pricing published by the corresponding providers (e.g., OpenAI, Claude, Fireworks AI, etc.). For unpopular models that are not served by commercial providers, we deploy them ourselves for experiments (only a few). In such cases, we approximate their costs using the pricing tiers published by commercial hosting platforms (e.g., Together.ai), which estimate serving costs based on model size (parameter count) and architecture type (e.g., MoE). Table 4 summarizes the pricing tiers we use for our self-hosted model.
(3) Routing Optimality.
This captures a router’s ability to perform optimal routing—that is, selecting the cheapest model that still produces a correct response. It consists of three sub-metrics: (a) Optimal Selection Ratio—the proportion of queries for which the router answers correctly by selecting the cheapest model; (b) Optimal Accuracy Ratio—the ratio between a router’s achieved accuracy and the upper-bound accuracy obtainable when always selecting the best model from its model pool; (c) Optimal Cost Ratio—the ratio between the cost incurred by the router’s selections and the cost of always choosing the optimal model. This metric will penalize routers that rely on unnecessarily expensive models when cheaper, correct alternatives are available. This metric will penalize routers that use unnecessarily expensive models when cheaper, correct alternatives are available.
(4) Routing Robustness.
This metric evaluates the router’s robustness against noisy inputs. We calculate it as the proportion of queries for which the router makes consistent routing decisions under perturbed input. Specifically, we generate noisy variants of queries—through paraphrasing, grammatical changes, synonym substitutions, and typos—and measure the percentage of cases where the router selects the same model as it does for the original, noise-free query. This captures the router’s capability for handling realistic, imperfect user queries.
(5) Routing Latency.
Since the router operates in the critical path of systems in production, it must handle millions of queries per second with minimal overhead. This metric measures the additional latency introduced by routing. It reflects the latency increase in both time-to-first-token (TTFT) and end-to-end response latency when a given router is employed.
5 RouterArena Evaluation Framework
5.1 Arena Ranking
RouterArena provides a series of router leaderboards that enables users to compare the capabilities of different routers and select the one best suited to their scenarios. It includes six ranking scores based on the evaluation metrics described in Section 4, including Arena, Optimal-selection-ratio, Optimal-acc-ratio, Optimal-cost-ratio, Robustness, and Latency. Among these, the Arena score captures the trade-off between accuracy and cost by combining them into a single composite measure using the Weighted Harmonic Mean [12]. Specifically, to better distinguish between routers with low costs, we apply a base-2 logarithmic ($log_{2}$) transformation to the cost values. Under this scaling, each doubling of price reduces the cost score by one unit. For router $i$ with cost $c_{i}$, we define its normalized cost as
$C_{i}=\frac{log_{2}(c_{max})-log_{2}c_{i}}{log_{2}(c_{max})-log_{2}(c_{min})}$
where $c_{\max}$ and $c_{\min}$ denote the maximum and minimum costs of routing 1k queries. Specifically, we choose $c_{min}=0.0044$, corresponding to the cost of the cheapest model in the leaderboard’s model pool. This reflects the cost of a router that always selects the cheapest model. We choose $c_{max}=200$ representing the most expansive model, OpenAI’s o1-Pro. This normalization maps the cost into range $[0,1]$, with larger values of $C_{i}$ corresponding to more economical routers. Next, we combine the normalized cost $C_{i}$ and accuracy $A_{i}$ using a weighted harmonic mean:
$S_{i,\beta}=\frac{(1+{\beta})A_{i}C_{i}}{{\beta}A_{i}+C_{i}},$
where the parameter $\beta>0$ controls the relative importance of accuracy versus cost. Setting $\beta>1$ places greater weight on cost. By default, we use $\beta=0.1$, emphasizing routing accuracy, because highly accurate routers are generally more valuable even if they incur slightly higher costs.
5.2 Automated Evaluation Framework
Although we demonstrate RouterArena with a specific set of routers in this paper, it is very easy to update the leaderboard with new routers. To facilitate this process, we have designed an automated evaluation framework that will be released publicly alongside the leaderboard. Figure 4 shows the overall system workflow. To evaluate a new router, the user can simply start by providing our framework with an access point (e.g., an API) to the router. The framework sends evaluation queries to the router, which performs routing inference on its end and returns its model selections. To ensure fairness, we run the inference ourselves and use cached results when possible, since many routers share overlapping model pools. During this process, the router’s response time is monitored to measure routing latency. Finally, the framework computes the evaluation metrics and aggregates the results, which are reflected in the leaderboard. Note that some commercial routers may not expose their model selections and instead return query answers directly. Such routers can still be evaluated with our framework, although certain metrics cannot be measured in this setting.
Figure 4: RouterArena Live Leaderboard.
6 Experiments
6.1 Experimental Settings
Router Selection.
For commercial routers, we evaluated the router from Not Diamond [25], which provides access to over 60 models, and the Azure Model Router [24], which currently only supports OpenAI models. We also included GPT-5 [27], whose model family incorporates an internal router. For NotDiamond, we selected 26 representative models spanning different parameter scales, architectures, and reasoning abilities. For Azure-Router, we evaluated the entire model pool, including GPT-5 model families. Appendix A provides full details of the model pools used for each router.
For open-source routers, we evaluated nine representative systems covering a diverse routing approaches. Specifically, we chose both the KNN- and MLP-based methods trained on RouterBench [15] as baselines. We further included GraphRouter [10], which leverages graph neural networks (GNNs) for routing, and the Universal Router [19], which uses K-means clustering. To capture cost–accuracy tradeoffs, we evaluated CARROT Router [31], while RouterDC [4] was incorporated as a dual contrastive learning–based approach. Additionally, we considered IRT-Router [32], which applies item response theory to explicitly model the interaction between query attributes and model capabilities, and RouteLLM [26], which performs binary selection between a stronger and a weaker model. Moreover, we also take the latest vLLM Semantic Router [35] into consideration, which leverages a ModernBERT [18] to categorize the incoming requests into pre-defined categories, and selects the model that has the highest score.
Router Training and Evaluation.
For commercial routers, no additional training is required; we simply accessed their provided APIs for evaluation. In contrast, for academia routers, we followed the training procedures and datasets specified in their open-source implementations. Specifically, we did not modify the training datasets or the task categorizations (if applicable). The model pools were configured in accordance with the original papers. In particular, for the vLLM-SR, we constructed the pool using both open-source models of varying parameter scales and proprietary models, with detailed configurations summarized in Table 3. After training each router, we evaluated them by feeding our benchmark dataset, recording the model selected, the latency incurred by the selection, and the confidence scores assigned to all candidate models in the pool.
6.2 Results
Figure 5: Deferral Curve: accuracy versus cost
Figure 6: Normalized Deferral Curve
Deferral Curve.
Figure 6 presents the trade-off between accuracy and inference cost. As we increase inference budget, we unlock more powerful models, driving the accuracy up. For open-source routers, we leveraged their confidence scores to apply budget-based masking, which produces multiple points along each curve. With only cheap models available, accuracy remains low, but as larger models enter the pool, routing accuracy increases. In contrast, commercial routers typically appear as single points because their model pools already include the best-performing models.
Two insights emerge. First, the orange dashed line shows the oracle accuracy, highlighting that all routers fall short of the best achievable performance. Second, the trade-off frontier differs by setting: commercial routers can achieve higher accuracy, but usually at significantly higher costs; open-source routers, on the other hand, achieve competitive performance at much lower budgets, though they plateau earlier. Notably, routers like CARROT and GraphRouter illustrate cost-efficient routing, while systems such as GPT-5 and NotDiamond lean heavily on expensive models for accuracy. This suggests that while commercial routers prioritize maximizing accuracy, academic approaches often explore the efficiency side of the frontier.
Normalized Deferral Curve.
Figure 6 reports router accuracy and cost normalized to each router’s best-performing model, point (100%, 100%) on the plot. The upper-left quadrant represents the ideal case—higher accuracy with lower cost by leveraging smaller models. In practice, most routers cluster near the baseline (100% cost, 100% accuracy), suggesting they over-rely on the strongest model and miss opportunities to defer to cheaper alternatives. Notably, NIRT-BERT illustrates inefficiency, reaching only baseline-level accuracy while incurring 378% of the cost.
By contrast, routers such as vLLM-SR and CARROT achieve meaningful savings: roughly 35% lower cost with under 2% accuracy degradation. These cases show routing can indeed improve efficiency when smaller models are effectively utilized. Overall, the figure highlights a clear trade-off—higher accuracy often comes with higher cost—while also pointing to promising directions for designing routers that move closer to the ideal frontier.
Figure 7: Actual and optimal accuracy, along with optimal selection ratio and cost ratio
Optimality Score.
Figure 7 highlights the inherent trade-off between routing accuracy and cost. In practice, routers that achieve higher accuracy typically do so at the expense of a higher cost ratio, since they defer more often to large, expensive models. This behavior lowers their optimal selection ratio, i.e., the frequency with which they choose the most efficient model for each query. This pattern is most apparent in the binary routers such as RouteLLM. By design, these routers face a sharp trade-off: they achieve higher accuracy by routing more queries to the stronger model, which drives up cost. In contrast, multi-model routers have a more flexible pool, and while the general trend still holds, we see greater variability depending on pool composition and routing strategy. Among non-binary routers, RouterDC stands out with the lowest cost ratio and highest optimal selection ratio, but this comes at the cost of poor overall accuracy. At the other extreme, MIRT-BERT achieves strong accuracy (close to 77% of its optimal accuracy) but requires nearly five times the optimal cost, placing it closer to the “high-cost high-accuracy” region of the trade-off frontier. In other words, while some routers are closer to the efficiency frontier than others, none simultaneously combine low cost and high accuracy. Overall, our findings indicate that current routing methods have learned to leverage large models to boost performance, but remain inefficient at recognizing when smaller models are sufficient. This creates a clear opportunity for future work: developing routers that can balance accuracy and efficiency by selectively deferring to large models only when necessary.
Figure 8: Router robustness and latency comparison.
Robustness and Latency.
Given that user prompts are often noisy, we further assess router sensitivity and robustness. Specifically, we prepend an irrelevant keyword to the input and observe whether the router alters its original model selection. We define robustness as 1 – the proportion of changed selections. As shown in Figure 8, routers leveraging latent representations of prompts demonstrate stronger stability against noise, whereas methods relying on explicit representations, such as KNN and NIRT-BERT, are considerably more sensitive. These findings highlight the importance of applying prompt engineering techniques to mitigate noisy queries in the future.
Furthermore, Figure 8 reports the end-to-end latency of routers on a single A100 GPU, measured from the time a request is received to the output of the model selection result. Among all methods, vLLM-SR and RouteLLM exhibit significantly higher latency because they rely on the OpenAI embedding API, which introduces additional network delays. In contrast, other routers consistently maintain sub-100ms latencies. Since the LLM router lies on the critical path of the end-to-end systems, our results provide new insights for industrial deployment: while LLM routers can optimize accuracy and cost, they also introduce non-negligible overhead that may even compromise service-level objectives (SLOs).
Table 2: Ranking of routers across multiple metrics. Lower values indicate better performance.
6.3 Insights from the Final RouterArena Leaderboard.
Our evaluation produces the router leaderboard shown in Table 2. The leaderboard consists of six ranking scores, and the overall ranking is determined by averaging across them. We highlight two key findings: (1) Commercial routers do not necessarily outperform open-source routers. For example, GPT-5 ranks #7 due to its restricted model pool, and NotDiamond ranks #12 because it frequently selects expensive models. (2) No router ranks at the top across all metrics, reflecting the inherent trade-offs in router design.
For developers and researchers, the findings highlight key deficiencies in current routing methods and point toward clear directions for designing the next generation of routers. The results show that all existing routers fall short of the oracle’s achievable performance, primarily because they are inefficient at recognizing when smaller, cheaper models are sufficient for a given query. Future work should focus on closing this performance gap. Moreover, the high latency and poor robustness of certain routers open new avenues of research beyond the traditional cost-versus-accuracy trade-off. Developers can use the platform’s automated framework to submit and benchmark new routers against established leaders, fostering innovation and transparently tracking progress in the field.
7 Related Work
LLM Router.
With the increasing availability of specialized models that can surpass even the most capable general-purpose LLMs in specific domains, both academia and industry have been actively exploring how to build LLM routers. In industry, several systems have emerged. Martian Router [37] proposed the idea of model mapping, while Storytell [33] categorizes user queries and routes them to the best-performing models. Other companies also seek to find the optimal model for user’s tasks by balancing performance and cost [25, 29, 27]. Recent academic efforts have also begun to emerge. GraphRouter [10] leverages graph neural networks, and Router-R1 [38] employs reinforcement learning. The growth of open-source solutions underscores the need for effective router evaluation [31, 32, 11].
LLM Router Benchmark.
RouterBench [15] introduces a large-scale dataset consisting of over 405k inference outcomes from representative LLMs. RouterEval [17] collects performance results from 8,500 LLMs across 12 widely used benchmarks. FusionBench [11] covers 14 tasks across five domains and leverages 20 open-source LLMs. Other benchmarks have also contributed to this line of work by using different data collection methods [20, 23]. However, these benchmarks fail to provide broad coverage across disciplinary domains or cover all kinds of routers. In contrast, our benchmark is systematically constructed based on an authoritative knowledge classification framework, making it the first comprehensive and actionable benchmark for LLM routing.
8 Conclusion
We introduce RouterArena, the first open platform for comprehensive router comparison. Our platform features a principled dataset with broad domain coverage and varying difficulty levels, an extensive set of evaluation metrics, and an automated framework to maintain a live leaderboard. Initial evaluations of 12 routers reveal a significant trade-off between accuracy and cost, showing that no single router is universally optimal. Commercial routers tend to achieve higher accuracy at a much greater expense, while open-source routers often present more cost-efficient solutions. Overall, our findings indicate that current routers are inefficient at leveraging cheaper models when appropriate, highlighting a clear opportunity for future work.
Appendix A Model Pools by Router
Table 3: Model pools used by different routers.
Router Model Pool
RouterBench WizardLM/WizardLM-13B-V1.2; claude-instant-v1; claude-v1; claude-v2; gpt-3.5-turbo-1106; gpt-4-1106-preview; meta/codellama-34b-instruct; meta/llama-2-70b-chat; mistralai/mistral-7b-chat; mistralai/mixtral-8x7b-chat; zero-one-ai/Yi-34B-Chat
GraphRouter meta-llama/llama-3-8b-instruct; mistralai/mixtral-8x7b-chat; nousresearch/nous-34b-chat; meta/llama-2-7b-chat; mistralai/mistral-7b-chat; meta/llama-3-70b-chat; meta/llama-3-turbo-8b-chat; meta/llama-3-turbo-70b-chat; meta/llama-3.1-turbo-70b-chat; qwen/qwen-1.5-72b-chat
Universal WizardLM/WizardLM-13B-V1.2; claude-instant-v1; claude-v1; claude-v2; gpt-3.5-turbo-1106; gpt-4-1106-preview; meta/codellama-34b-instruct; meta/llama-2-70b-chat; mistralai/mistral-7b-chat; mistralai/mixtral-8x7b-chat; zero-one-ai/Yi-34B-Chat
CarrotRouter aws-claude-3-5-sonnet-v1; aws-titan-text-premier-v1; openai-gpt-4o; openai-gpt-4o-mini; wxai-granite-3-2b-instruct-8k-max-tokens; wxai-granite-3-8b-instruct-8k-max-tokens; wxai-llama-3-1-70b-instruct; wxai-llama-3-1-8b-instruct; wxai-llama-3-2-1b-instruct; wxai-llama-3-2-3b-instruct; wxai-llama-3-3-70b-instruct; wxai-mixtral-8x7b-instruct-v01; wxai-llama-3-405b-instruct
RouterDC mistralai/Mistral-7B-v0.1; meta-math/MetaMath-Mistral-7B; itpossible/Chinese-Mistral-7B-v0.1; HuggingFaceH4/zephyr-7b-beta; cognitivecomputations/dolphin-2.6-mistral-7b; meta-llama/llama-3-8b-instruct; cognitivecomputations/dolphin-2.9-llama3-8b
IRT-Router glm_4_air; glm_4_flash; glm_4_plus; gpt_4o; gpt_4o_mini; gpt_4o_mini_cot; deepseek_coder; deepseek_chat; qwen25_32b_int4; qwen25_7b_instruct; qwen25_72b_instruct; qwq_32b_preview; qwen25_math_7b_instruct; llama31_8b_instruct; llama31_70b_instruct; llama31_405b_instruct; mixtral_8x7b_instruct; mistral_7b_instruct_v02; ministral_8b_instruct_2410; gemini15_flash; claude35_haiku20241022
RouteLLM openai-gpt-4o; mixtral_8x7b_instruct
We provide the model pool used by each router in Table 3. Each router is trained on its own distinct set of LLMs. When replicating training runs, the same LLM allocation is preserved for each router to maintain consistency in usage. Since some models do not have a serverless inference endpoint, we utilized vLLM [21] to perform inference on them using an 8x A100 40 GB server. The Together.ai price bracket in Fig 4 estimates the inference price based on the parameter size.
Table 4: Dense and MoE large language model size and price per million tokens.
Dense Models Mixture-of-Experts Models
Model Size Price Model Size Price
Up to 4B \$0.10 Up to 56B \$0.60
4.1B – 8B \$0.20 56.1B – 176B \$1.20
8.1B – 21B \$0.30 176.1B – 480B \$2.40
21.1B – 41B \$0.80
41.1B – 80B \$0.90
80.1B – 110B \$1.80
Appendix B RouterArena Dataset
B.1 Datasets Details
Figure 9: The composition datasets in RouterArena.
Figure 9 illustrates the domain coverage of our dataset across the nine Dewey Decimal categories. Each horizontal bar represents the relative contribution of different source datasets within a category. This distribution highlights the systematic integration of multiple datasets to achieve a more balanced representation of both general-purpose and highly specialized domains.
Note, we opt out of the domain of Religion due to its sensitivity in its content of the topic.
Table 5: Overview of dataset columns
Column Description Example
Domain Bloom’s taxonomy high-level class 9 History
Category Bloom’s taxonomy sub-class 02 Library and information sciences
Dataset Name Source dataset ArcMMLU
Global Index Unique instance ID ArcMMLU_114
Context Supporting passage (if any) Sasha decided to watch TV and get some food
Question Input question What is the capital of France?
Options Multiple-choice options [“10”, “20”, “30”, “40”]
Answer Ground-truth answer Paris
Bloom Level Bloom’s taxonomy difficulty level Understanding
We provide detailed dataset columns in Table 5. Domain and Category represent the higher and lower classification levels defined by Dewey Decimal Classification (DDC) [8]. Dataset Name represents the specific dataset that the question comes from. Global Index is a unique identifier for each question across the benchmark dataset. Context, Question, and Options are question-specific fields, and may be used to provide more information to curate the input question after formatting it into the input prompt. Answer is the golden answer. Bloom Level is the classified difficulty level based on Bloom’s Taxonomy [1].
Figure 10: The difficulty distribution in RouterArena.
Figure 10 shows the overall difficulty distribution in our benchmark dataset.
B.2 Dataset Examples
Figure 11: Dataset examples.
Figure 11 shows some examples from the benchmark dataset.
Appendix C Leaderboard
Figure 12 presents the spider plot of RouterArena, which compares six routing methods (CARROT, RouterDC, GraphRouter, MIRT-BERT, NIRT-BERT, and RouteLLM) across five evaluation dimensions: Arena Score, Cost-ratio Score, Optimal-acc Score, Latency Score, and Robustness Score. Each axis indicates higher performance in the outward direction, allowing a direct visualization of trade-offs. For example, CARROT achieves strong performance in Arena and Latency Scores, while RouterDC excels in Cost-ratio Score.
Figure 12: The spider plot of RouterArena
Appendix D LLM Prompts
D.1 Bloom Taxonomy Prompt
We used the following prompt for classifying each question into different Bloom Levels:
Instruction: You are an evaluator tasked with classifying questions by cognitive difficulty using Bloom’s Taxonomy (Revised 2001). Bloom’s Taxonomy defines six levels of cognitive processes: 1. Remember – Recall or recognize facts, terms, or concepts. 2. Understand – Explain, summarize, interpret, or demonstrate comprehension. 3. Apply – Use learned knowledge to solve problems in new or routine situations. 4. Analyze – Break down information, examine parts, relationships, or underlying causes. 5. Evaluate – Make judgments or decisions based on evidence, criteria, or standards. 6. Create – Put elements together to form a new structure, idea, or product. Your Task: Given the details of a question, determine which Bloom’s level best represents the cognitive process required to answer it. - Output the result in a structured JSON block with no additional text. - If multiple levels might apply, choose the highest level required. Output Format:
  "bloom_level": "<Remember | Understand | Apply | Analyze |
  Evaluate | Create>"
Input Fields:
- Dataset name: {dataset_name}
- Dataset classification (DDC category): {dataset_category}
- Question: {question}
- Question Index: {question_index}
- Context (if any): {context}
- Options (if MCQ): {options}
- Answer: {answer}
D.2 Evaluation Prompt
The following is an example of an evaluation prompt we use:
Please read the following multiple-choice questions and
[…]
${Context}\n\n$
${Question}\n\n$
$\n{Options}\n\n$
[…]
the correct letter choice.
Keep the explanation or feedback within 3 sentences.